OpenStudy (anonymous):

Integrate:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 5-\sqrt{x} }{ 5+\sqrt{x} }\]

OpenStudy (aravindg):

dx

OpenStudy (anonymous):

yeah, sorry, dx

OpenStudy (anonymous):

Well, here's as far as I've gotten...

OpenStudy (anonymous):

substitute sqrt(x) for u

OpenStudy (anonymous):

I end up with \[2\int\limits_{}^{}\frac{ 5u-u ^{2} }{ u+5 }du\]

OpenStudy (anonymous):

Then I do long division.

OpenStudy (anonymous):

I'd first rationalise the denominator of the integrand: \[\frac{5-\sqrt{x}}{5+\sqrt{x}}\frac{5-\sqrt{x}}{5-\sqrt{x}}=\frac{25-10\sqrt{x}+x}{25-x}\]which then splits our integral to three distinct ones :\[25\int_{}{}\frac{dx}{25-x}\](substitute and you''ll get a natural log ),\[-10\int_{}{}\frac{\sqrt{x}}{25-x}dx\]( suggestions ? ),\[\int_{}{}\frac{x}{25-x}dx\](integration by parts).

OpenStudy (anonymous):

\[2\int\limits_{}^{}10-u-\frac{ 50 }{ u+5 }du\]

OpenStudy (anonymous):

I tried what you did first, then got stuck with the second part of the integral.

OpenStudy (anonymous):

So I got \[20\sqrt{x}-x-100\ln \left| \sqrt{x}+5 \right|+C\]

OpenStudy (anonymous):

But it looks wrong to me...

OpenStudy (anonymous):

One way to check is to take the derivative of your answer ..... :-)

OpenStudy (anonymous):

Hmm... ok. I'll try that.

OpenStudy (anonymous):

Man, trying to differentiate this and place it back in that form is as hard as integrating it. Haha.

OpenStudy (anonymous):

Yeah, I can't make it match. I must have done something wrong.

OpenStudy (anonymous):

I'll check you integration ....

OpenStudy (anonymous):

OK\[u=\sqrt{x}=x^{1/2}\]gives\[\frac{du}{dx}=(1/2)x^{-1/2}=\frac{1}{2\sqrt{x}}\]thus\[dx=2\sqrt{x}du=2u\times du\]and\[I= \int_{}{}\frac{5-\sqrt{x}}{5+\sqrt{x}}dx=2\int_{}{}\frac{(5-u)u}{5+u}du\]now\[\frac{-u^2 +5u}{u+5}=\frac{-u^2-5u+5u+5u}{u+5}\]\[=\frac{-u(u+5)+10u}{u+5}=-u + \frac{10u}{u+5}\]\[=-u +\frac{10u+50-50}{u+5}=-u+\frac{10(u+5)-50}{u+5}\]\[=-u + 10-\frac{50}{u+5}\]so\[I=2\int_{}{}[-u + 10-\frac{50}{u+5}]du\]\[=2\frac{-u^{2}}{2}+20u-100ln|u+5|+C\]\[=-x+20\sqrt{x}-100ln|5+\sqrt{x}|+C\]well, wise men agree and fools never differ ! :-)

OpenStudy (anonymous):

:) So, does that mean this is the correct answer?

OpenStudy (anonymous):

Yeah, I'd pay it. :-)